OpenStudy (erniellyd28):

find the slope intercept form of the line whose slope is 3 and that passes throught the point (-5,8)

OpenStudy (roxy.girl):

SO you would make a graph?

OpenStudy (mww):

The slope intercept form of a line is y = mx + b where m is your gradient and b is the y intercept. So sub in your slope and point into this equation. This will allow you to find the value of b, your y-intercept. Then resub this b into y = mx +b.

OpenStudy (wolf1728):

The equation has the general form y=mx +b b is calculated by this equation: b = y -mx Use the point (x= -5, y =8) and the slope "m" equals 3 That should help

OpenStudy (roxy.girl):

yes that is right! :)